OpenStudy (anonymous):

[8.06] Select the factors of x2 − 14x + 49. (x + 49)(x + 1) (x + 7)(x + 7) (x − 7)(x − 7) (x − 49)(x − 1)

OpenStudy (anonymous):

you are supposed to recognize this as a perfect square, since \(49=7^2\) and \(2\times 7=14\)

OpenStudy (anonymous):

it is the square \((x-7)^2\) even though that is not one of your answer choices try instead \((x-7)(x-7)\)

OpenStudy (anonymous):

so its C? right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is almost always C anyway